### v1

@@ Appendix C Evaluation Metrics
  • Pass@k (Kulal et al., 2019; Chen et al., 2021): We use the conventional Pass@k, which measures the fraction of problems for which at least one of the $k$ generated solutions is correct. We use $k=8$.
@@
  • Olympics Medal System: It uses the authoritative cutoffs in the Olympiads to decide if a model’s performance is qualified for a medal (gold, silver, or bronze).
@@ Appendix D Full Results
We present the complete evaluation results of all models on LiveOIBench. Table A9 provides the overall leaderboard across all $72$ contests, while Table A10 breaks down performance by contest tags. Finally, Figure A1 shows a screenshot of the LiveOIBench website, which allows users to interactively explore model performances by selecting specific contest ranges.
Model Medals Relative Human Pass Rate (%) Elo Rating Division Elo Rating
Gold(%) Silver(%) Bronze(%) Medals(%) Score(%) Percentile D1 D2 D3 D4
@@
GPT-5 50.00 30.56 8.33 88.89 67.21 81.76 63.03 2414 2426 2322 2412 2583
Gemini-2.5-Pro 31.94 22.22 23.61 77.78 51.33 71.80 44.46 2192 1963 2028 2308 2551
GPT-O3-Mini-High 26.39 23.61 22.22 72.22 47.69 64.28 44.19 2088 1807 1894 2284 2449
Gemini-2.5-Flash 15.28 23.61 23.61 62.5 41.29 56.81 36.06 1945 1700 1700 2091 2505
GPT-4.1 4.17 13.89 22.22 40.28 24.78 35.99 18.32 1482 1339 1134 1724 1994
@@
GPT-OSS-120B-High 50.00 26.39 11.11 87.50 62.78 72.88 60.14 2205 1950 2122 2264 2520
GPT-OSS-20B-High 22.22 29.17 23.61 75.00 49.55 57.72 52.81 2020 1763 1797 2167 2504
GPT-OSS-120B 29.17 23.61 20.83 73.61 49.23 59.90 47.78 2032 1638 1894 2193 2493
GPT-OSS-20B 19.44 23.61 25.00 68.06 42.36 53.94 42.80 1901 1501 1660 2165 2383
Qwen3-32B 9.72 15.28 29.17 54.17 32.86 42.00 27.70 1665 1342 1455 1959 2022
DeepSeek-R1 6.94 19.44 26.39 52.78 33.43 42.29 28.87 1617 1443 1278 1906 2015
Qwen3-14B 5.56 15.28 25.0 45.83 27.24 34.59 22.73 1402 976 1241 1652 1938
QWQ-32B 5.56 13.89 26.39 45.83 26.56 33.84 23.95 1491 1281 1113 1877 1956
Qwen3-30B 5.56 20.83 18.06 44.44 27.68 36.69 23.18 1549 1201 1323 1862 1995
Qwen3-8B 1.39 12.5 26.39 40.28 24.25 31.03 19.05 1426 1206 1312 1534 1789
DeepSeek-R1-Distill-Llama-70B 1.39 8.33 23.61 33.33 20.50 32.30 16.88 1283 1042 1103 1472 1665
DeepSeek-R1-Distill-Qwen-32B 1.39 8.33 20.83 30.56 19.14 27.03 14.86 1284 964 1074 1631 1549
Qwen3-4B 1.39 8.33 16.67 26.39 16.81 24.28 13.61 1153 970 897 1332 1622
DeepSeek-R1-Distill-Qwen-14B 1.39 2.78 9.72 13.89 13.41 22.77 10.56 1089 897 991 1166 1457
DeepSeek-R1-Distill-Llama-8B 0.0 0.0 2.78 2.78 3.10 11.86 2.46 724 724 628 705 1103
@@
DeepSeek-V3 4.17 8.33 22.22 34.72 21.70 31.76 17.10 1283 1239 1187 1598 1827
Qwen3-32B-Non-Thinking 1.39 4.17 11.11 16.67 12.92 24.64 8.78 1040 957 844 1227 1251
Qwen2.5-Coder-32B-Instruct 1.39 2.78 9.72 13.89 11.25 19.90 6.15 1023 983 701 1247 1384
Qwen2.5-Coder-14B-Instruct 1.39 2.78 6.94 11.11 9.66 19.56 5.53 966 935 849 969 1360
Mistral-Large-Instruct-2411 1.39 1.39 8.33 11.11 9.99 18.70 5.90 1023 939 875 1122 1376
Mistral-Small-3.1-24B-2503 1.39 0.0 9.72 11.11 7.75 19.08 4.75 909 805 822 879 1334
Llama-4-Scout 1.39 1.39 5.56 8.33 9.88 19.60 6.32 1008 825 892 1107 1316
Qwen2.5-72B 1.39 2.78 5.56 9.72 9.90 19.24 5.55 1000 875 862 1022 1508
Llama-3.3-70B-Instruct 0.0 1.39 8.33 9.72 10.00 21.37 5.65 1056 899 1069 1020 1458
Qwen3-30B-Non-Thinking 1.39 0.0 6.94 8.33 10.48 17.28 6.99 989 962 791 1052 1425
Qwen3-4B-Non-Thinking 0.0 1.39 5.56 6.94 6.65 15.30 4.47 894 818 753 932 1303
Qwen3-8B-Non-Thinking 0.0 1.39 2.78 4.17 7.53 16.82 4.04 843 745 701 842 1357
CODESTRAL-22B-V0.1 0.0 1.39 2.78 4.17 6.84 15.94 4.34 912 948 784 895 1275
Llama-3.1-8B-Instruct 0.0 1.39 1.39 2.78 4.19 13.49 2.45 761 714 644 808 1073
Table A9: Main results of all models we have evaluated on all 72 contests from LiveOIBench.
@@
Figure A1: LiveOIBench website that displays leaderboard across models
@@ E.2 Inference-Time Scaling
Inference-time scaling has been shown effective for improving model performance in math (Snell et al., 2024; Brown et al., 2024) and coding (Li et al., 2025a; Ehrlich et al., 2025) domains. We investigate two dimensions: parallel scaling involves sampling multiple diverse solution candidates (Chen et al., 2021; Jain et al., 2024), while sequential scaling generates long chains-of-thought with complex reasoning strategies such as self-reflection and backtracking (DeepSeek-AI et al., 2025).
@@
Parallel Scaling: GPT-5 demonstrates superior coding capacity boundary. Figure 2 reveals significant differences in coding capacity boundaries (Yue et al., 2025) across models as measured by Pass@k. GPT-5 could pass around $64\%$ of the problems given 8 attempts per problem. The steepest improvements occur between Pass@1 and Pass@4, indicating that the marginal benefit of additional attempts diminishes rapidly as models approach their capacity limits (Kulal et al., 2019). The persistent performance gaps between proprietary and open-source models across all sampling levels suggest fundamental differences in maximum coding capability rather than artifacts of insufficient attempts (Li et al., 2022b; Hendrycks et al., 2021a).
@@
Both scaling approaches provide complementary benefits but face efficiency limitations. Sequential scaling shows promise for complex algorithmic problems but requires substantial computational resources, while parallel scaling reveals each model’s performance ceiling as improvements plateau with additional samples (Chen et al., 2021; Austin et al., 2021). Future work could focus on developing hybrid approaches that combine both scaling paradigms while reducing computational overhead.
E.3 Reasoning Behaviors Analysis

### v2

@@ Appendix C Evaluation Metrics
  • Pass@k (passk1; chen2021codex): We use the conventional Pass@k, which measures the fraction of problems for which at least one of the $k$ generated solutions is correct. We use $k=8$.
@@
  • Olympiad Medal System: It uses the authoritative cutoffs in the Olympiads to decide if a model’s performance is qualified for a medal (gold, silver, or bronze).
@@ Appendix D Full Results
We present the complete evaluation results of all models on LiveOIBench. Table A9 provides the overall leaderboard across all $72$ contests, while Table A10 breaks down performance by contest tags.
Model Medals (%) Relative Human Percentile (%) Pass Rate (%) CF Rating
Gold Silver Bronze Total Score (%) All D1 D2 D3 D4 All Easy Med. Hard All D1 D2 D3 D4
@@
GPT-5 50.00 30.56 8.33 88.89 67.21 81.76 73.68 79.30 84.87 97.20 63.03 92.25 60.42 34.19 2414 2426 2322 2412 2583
Grok-4-Fast-Reasoning 45.83 20.83 16.67 83.33 56.99 74.23 63.79 67.62 79.14 97.46 50.95 70.42 39.58 15.38 2221 2053 2158 2224 2598
Gemini-2.5-Pro 31.94 22.22 23.61 77.78 51.33 71.80 55.84 65.23 79.02 95.90 44.46 82.39 32.64 16.24 2192 1963 2028 2308 2551
GPT-O3-Mini-High 26.39 23.61 22.22 72.22 47.69 64.28 53.07 50.76 75.79 94.67 44.19 84.51 34.03 11.11 2088 1807 1894 2284 2449
Gemini-2.5-Flash 15.28 23.61 23.61 62.50 41.29 56.81 43.48 43.79 69.89 93.97 36.06 75.35 22.92 6.84 1945 1700 1700 2091 2505
Claude-Sonnet-4.5 11.11 20.83 34.72 66.67 38.30 53.08 36.69 49.80 65.79 85.63 27.05 61.97 9.72 5.13 1848 1766 1595 2057 2060
GPT-4.1 4.17 13.89 22.22 40.28 24.78 35.99 21.47 24.90 47.56 79.73 18.32 48.59 4.86 3.42 1482 1339 1134 1724 1994
@@
GPT-OSS-120B-High 50.00 26.39 11.11 87.50 62.78 72.88 69.46 63.36 80.92 96.82 60.14 92.25 54.86 30.77 2205 1950 2122 2264 2520
GPT-OSS-20B-High 22.22 29.17 23.61 75.00 49.55 57.72 47.18 44.47 72.55 94.07 52.81 86.62 42.36 17.95 2020 1763 1797 2167 2504
GPT-OSS-120B 29.17 23.61 20.83 73.61 49.23 59.90 53.89 43.75 72.87 91.14 47.78 87.32 33.33 19.66 2032 1638 1894 2193 2493
GPT-OSS-20B 19.44 23.61 25.00 68.06 42.36 53.94 48.43 37.94 66.58 88.26 42.80 81.69 29.86 12.82 1901 1501 1660 2165 2383
Qwen3-32B 9.72 15.28 29.17 54.17 32.86 42.00 31.36 31.17 56.64 81.73 27.70 67.61 6.25 6.84 1665 1342 1455 1959 2022
DeepSeek-R1 6.94 19.44 26.39 52.78 33.43 42.29 30.22 27.05 55.89 80.92 28.87 69.01 8.33 6.84 1617 1443 1278 1906 2015
Qwen3-14B 5.56 15.28 25.00 45.83 27.24 34.59 24.51 25.66 47.57 74.78 22.73 58.45 5.90 7.69 1402 976 1241 1652 1938
QWQ-32B 5.56 13.89 26.39 45.83 26.56 33.84 19.15 26.44 49.93 71.94 23.95 57.75 18.31 11.97 1491 1281 1113 1877 1956
Qwen3-30B 5.56 20.83 18.06 44.44 27.68 36.69 24.94 27.25 47.33 72.15 23.18 63.38 14.93 7.26 1549 1201 1323 1862 1995
Qwen3-8B 1.39 12.50 26.39 40.28 24.25 31.03 25.37 23.23 40.68 69.90 19.05 53.52 15.56 7.26 1426 1206 1312 1534 1789
DeepSeek-R1-Distill-Llama-70B 1.39 8.33 23.61 33.33 20.50 32.30 22.28 23.63 39.29 76.89 16.88 43.66 15.85 7.69 1283 1042 1103 1472 1665
DeepSeek-R1-Distill-Qwen-32B 1.39 8.33 20.83 30.56 19.14 27.03 11.43 15.59 37.55 44.75 14.86 40.14 1.39 1.71 1284 964 1074 1631 1549
Qwen3-4B 1.39 8.33 16.67 26.39 16.81 24.28 18.35 16.17 29.91 54.21 13.61 31.08 12.60 5.13 1153 970 897 1332 1622
DeepSeek-R1-Distill-Qwen-14B 1.39 2.78 9.72 13.89 13.41 22.77 15.33 15.23 25.64 47.24 10.56 25.35 9.15 4.05 1089 897 991 1166 1457
DeepSeek-R1-Distill-Llama-8B 0.00 0.00 2.78 2.78 3.10 11.86 10.82 10.37 9.51 33.46 2.46 10.56 1.56 0.00 724 724 628 705 1103
@@
DeepSeek-V3 4.17 8.33 22.22 34.72 21.70 31.76 16.42 20.86 41.88 73.73 17.10 43.66 2.78 2.56 1283 1239 1187 1598 1827
Qwen3-32B-Non-Thinking 1.39 4.17 11.11 16.67 12.92 24.64 14.14 17.35 24.10 52.37 8.78 21.13 0.69 4.27 1040 957 844 1227 1251
Qwen2.5-Coder-32B-Instruct 1.39 2.78 9.72 13.89 11.25 19.90 15.31 17.26 22.73 48.38 6.15 22.54 5.90 1.71 1023 983 701 1247 1384
Qwen2.5-Coder-14B-Instruct 1.39 2.78 6.94 11.11 9.66 19.56 13.48 14.56 23.55 46.20 5.53 20.42 4.51 1.71 966 935 849 969 1360
Mistral-Large-Instruct-2411 1.39 1.39 8.33 11.11 9.99 18.70 14.62 15.13 22.66 43.03 5.90 21.13 4.65 1.71 1023 939 875 1122 1376
Mistral-Small-3.1-24B-2503 1.39 0.00 9.72 11.11 7.75 19.08 10.67 10.92 20.04 42.86 4.75 17.61 3.41 1.71 909 805 822 879 1334
Llama-4-Scout 1.39 1.39 5.56 8.33 9.88 19.60 10.67 11.97 24.11 47.09 6.32 21.83 4.34 1.71 1008 825 892 1107 1316
Qwen2.5-72B 1.39 2.78 5.56 9.72 9.90 19.24 10.84 12.99 22.96 45.43 5.55 14.08 0.00 1.71 1000 875 862 1022 1508
Llama-3.3-70B-Instruct 0.00 1.39 8.33 9.72 10.00 21.37 15.04 16.67 26.37 50.32 5.65 20.42 4.79 1.71 1056 899 1069 1020 1458
Qwen3-30B-Non-Thinking 1.39 0.00 6.94 8.33 10.48 17.28 10.60 12.11 22.83 43.92 6.99 23.94 5.49 1.71 989 962 791 1052 1425
Qwen3-4B-Non-Thinking 0.00 1.39 5.56 6.94 6.65 15.30 6.65 8.76 10.91 44.36 4.47 10.56 0.00 2.56 894 818 753 932 1303
Qwen3-8B-Non-Thinking 0.00 1.39 2.78 4.17 7.53 16.82 9.78 9.01 14.16 39.65 4.04 10.56 0.00 1.71 843 745 701 842 1357
CODESTRAL-22B-V0.1 0.00 1.39 2.78 4.17 6.84 15.94 9.81 7.46 6.86 42.39 4.34 8.45 1.39 2.56 912 948 784 895 1275
Llama-3.1-8B-Instruct 0.00 1.39 1.39 2.78 4.19 13.49 7.90 6.15 7.60 30.13 2.45 5.63 0.00 0.85 761 714 644 808 1073
Table A9: Main results of all models evaluated on 72 contests, with expanded Human Percentile, Pass Rate, and CF rating breakdowns.
@@ E.2 Inference-Time Scaling
Inference-time scaling has been shown effective for improving model performance in math (DBLP:journals/corr/abs-2408-03314; DBLP:journals/corr/abs-2407-21787) and coding (DBLP:journals/corr/abs-2502-14382; DBLP:journals/corr/abs-2501-14723) domains. We investigate two dimensions: parallel scaling involves sampling multiple diverse solution candidates (chen2021codex; jain2024livecodebenchholisticcontaminationfree), while sequential scaling generates long chains-of-thought with complex reasoning strategies such as self-reflection and backtracking (deepseek-r1).
@@
Parallel Scaling: GPT-5 demonstrates superior coding capacity boundary. Figure 2 reveals significant differences in coding capacity boundaries (DBLP:journals/corr/abs-2504-13837) across models as measured by Pass@k. GPT-5 could pass around $64\%$ of the problems given 8 attempts per problem. The steepest improvements occur between Pass@1 and Pass@4, indicating that the marginal benefit of additional attempts diminishes rapidly as models approach their capacity limits (passk1). The persistent performance gaps between proprietary and open-source models across all sampling levels suggest fundamental differences in maximum coding capability rather than artifacts of insufficient attempts (Li2022; Hendrycks2021).
@@
Both scaling approaches provide complementary benefits but face efficiency limitations. Sequential scaling shows promise for complex algorithmic problems but requires substantial computational resources, while parallel scaling reveals each model’s performance ceiling as improvements plateau with additional samples (chen2021codex; austin2021programsynthesislargelanguage). Future work could focus on developing hybrid approaches that combine both scaling paradigms while reducing computational overhead.
Model Relative Score (%) Pass Rate (%) Gold (%) Silver (%) Bronze (%) Total Medals (%)
GPT-OSS-120B_Python 44.35 46.97 13.64 13.64 27.27 54.55
GPT-OSS-120B_Java 50.78 56.82 9.09 36.36 27.27 72.73
GPT-OSS-120B_C++ 53.26 59.09 27.27 31.82 13.64 72.73
Table A11: Performance comparison of GPT-OSS-120B across Python, Java, and C++ on 22 USACO competitive programming contests. Metrics shown include relative score, pass rate, and medal distributions (gold, silver, bronze, and total medal percentage). Results indicate that C++ achieves the highest overall performance, followed by Java, with Python exhibiting significantly lower performance, likely due to differences in execution speed, memory management, and efficiency of standard libraries.
E.3 Performance Variations Across Programming Languages
To provide deeper insights into language-specific model capabilities, we present a comparative analysis of GPT-OSS-120B’s performance across Python, Java, and C++, evaluated on a subset of 132 problems from the USACO contest series. As shown in the Table A11, the model performs best when generating solutions in C++, highlighting the inherent advantage of C++ in competitive programming due to its superior execution speed, precise memory control, and optimized functions. Java delivers intermediate performance; although robust and widely supported, it incurs overhead from JVM execution and automated memory management, placing it slightly behind C++. Meanwhile, although Python is the most resourceful and extensively studied language for evaluating LLM coding capability, it shows a considerable performance gap due to its interpreted nature and dynamic typing, sacrificing critical speed and efficiency. Thus, contest results clearly show C++ performing best, followed by Java, with Python significantly behind.
Figure A4: Histogram showing reasoning behaviors across Python, Java, and C++ are similar, which suggests the performance gain of using C++ mainly comes from the language effiency.
To better understand the reasons behind these performance differences, we further analyzed GPT-OSS-120B’s reasoning behaviors across these programming languages. As illustrated in the Figure A4, the model consistently allocates reasoning tokens similarly across Python, Java, and C++, indicating that its reasoning strategy remains stable irrespective of language choice. This stability in reasoning behavior suggests that the performance advantage observed for C++ primarily stems from computational efficiency rather than differences in the underlying reasoning approach.
E.4 Impact of Template Similarity on Model Performance
To investigate if models exhibit similar performance on problems sharing similar solution templates, we compute the official solution similarity for each problem pair and analyze its relationship with the models’ performance differences on these problems. Specifically, we assessed GPT-5 and Grok-4-fast-reasoning by plotting the absolute difference in their relative scores against each problem pair’s solution similarity (Figures A5). If template-based contamination significantly influenced model performance, we would expect to observe consistently small performance differences between problem pairs with similar official solutions. However, our analysis reveals only a very weak and negligible correlation (GPT-5: Pearson’s r = -0.084; Grok-4: Pearson’s r = -0.072). This indicates that even problem pairs with similar official solution structures often result in substantially different model performances. Additionally, we found that most problem pairs exhibit low similarity, emphasizing the wide variety of problem-solving approaches necessary for success on our benchmark.
Figure A5: Scatter plots of solution similarity vs. model performance differences for GPT-5 (left) and Grok-4 (right). Weak correlations (GPT-5: r=0.08r=-0.08; Grok-4: r=0.07r=-0.07) indicate minimal template-based contamination.
Figure A6: No significant positive correlation is observed between GPT-OSS-120B’s familiarity with task statements and solutions (normalized via min-max scaling) and its performance, indicating that higher familiarity does not necessarily translate to better outcomes.
Figure A7: Histogram showing similarity scores between accepted model solutions (GPT-5 and Grok-4-Fast-Reasoning) and their corresponding official solutions. All similarity scores remain below 50%, indicating that accepted model solutions are substantially distinct from official solutions. Density represents the probability per unit similarity, so that the total area of the histogram equals 1.
Model Δ (%) IM MA AH PS SO GR GTR BS NT GT DS CB DP TR ST Avg
[…]
Table A12: Percentage improvements across algorithmic task categories from scaling model parameters and reasoning-token budgets.
E.5 More analysis on how different attributes can impact the model performance
In Table A10 and Table A12, we present a detailed analysis of model performance across various algorithmic categories, examining three primary factors: model size, reasoning-token budget, and training strategies (RL and distillation).
Model Size: Scaling GPT-OSS from 20B to 120B parameters under low-reasoning effort yields an average improvement of +18%. This gain is concentrated mainly in Greedy (+83.30%), Graph Traversal (+42.80%), Combinatorics (+33.40%), and Tree (+33.40%). These patterns indicate that additional parameters enhance the model’s structural representations and pattern-recognition capacity, enabling better handling of connectivity and global relational structure.
Reasoning Tokens: Increasing the reasoning-token budget from low to high, while maintaining the same GPT-OSS-20B model weights, produces a substantially larger average improvement of +113%. Gains are most pronounced in algorithmic tasks such as string manipulation (+268.33%), tree traversal (+220.00%), greedy algorithms (+212.64%), dynamic programming (+186.43%), and sorting (+121.44%), all of which demand extended multi-step, sequential reasoning. Notably, greedy and tree traversal appear in the top-5 for both factors: larger models recognize their structural patterns more reliably, while deeper reasoning is essential to execute the multi-step, stateful chains these problems demand.
RL vs. Distillation: Since each model family differs in architecture, pretraining data, and training pipelines, it’s difficult to cleanly isolate the effects of RL versus distillation on downstream performance. However, preliminary comparisons between Qwen3-14B (a distilled model) and GPT-OSS-20B (trained via SFT and RL) suggest RL-trained models, with SFT warmup and enough RL training, can substantially outperform distilled models, with an average improvement of +118%. Isolating the contribution of each training stage is an important future direction for us to understand how these methods shape algorithmic skills.
Overall, our analyses show that reasoning token allocation, far more than model size, drives improvements on the most algorithmically complex tasks. Also, directly trained RL models perform better than distilled models.
Model Relative Score (%) Pass Rate (%) Human Percentile
GPT-OSS-120B-P1 47.76 45.85 58.50
GPT-OSS-120B-P2 48.07 47.10 59.76
GPT-OSS-120B-P3 46.75 45.69 57.73
GPT-OSS-120B-P4 48.04 45.68 59.49
GPT-OSS-120B-Medium 48.88 47.33 58.10
Mean ± Std 47.90 ± 0.69 46.33 ± 0.73 58.72 ± 0.79
Table A13: Performance of GPT-OSS-120B-Medium under different prompt variants.
E.6 Prompt Sensitivity
We generate four alternative prompt variants using GPT-5 and run GPT-OSS-120B-Medium with four additional prompt variants to explore how different instructions might influence the model’s performance. While these prompts do not represent all possible instruction styles, we observe minimal performance differences across them, suggesting that prompt variants have little influence on model performance. The default prompt and prompt variants can be found at B and E.7.
In Table A13, across all five prompt variants, GPT-OSS-120B achieves $47.9\pm 0.7\%$ relative score and $46.3\pm 0.7\%$ pass rate, with human percentile tightly clustered around $58.7\pm 0.8$, indicating low sensitivity to prompt choices. A more systematic and large-scale study of prompt sensitivity is an interesting direction for future work.
E.7 Analysis of Template-Based Contamination and Model Performance
We selected CodeContest (alphacode; wang2025codecontestshighqualitytestcase) as our seed dataset to represent template solutions commonly found in training data, as it is among the most prominent competitive coding datasets with over 10,000 problems. For each problem, we sampled three C++ solutions, filtering out solutions shorter than 20 lines or longer than 1,000 lines. Using Dolos, we calculated the similarity between official solutions and these template solutions, recording the highest similarity score for each official solution. A higher similarity score indicates a greater likelihood of template-based contamination.
To examine whether template similarity influences model performance, we conducted correlation analyses using GPT-5 and Grok-4-Fast-Reasoning. In Figure A8, our analysis revealed weak correlations between template similarity and relative performance scores: GPT-5 exhibited a Pearson correlation coefficient of 0.136, and Grok-4 demonstrated a Pearson correlation coefficient of 0.193. These correlations, though slightly stronger than those observed in our previous analysis, remain limited.
This finding indicates that template similarity contributes minimally to model performance. Rather, the dominant factor influencing performance is the model’s intrinsic problem-solving and reasoning capabilities. Given the nature of competitive coding problems, models rely on learned algorithmic strategies and analytical techniques to solve problems effectively. Therefore, we contend that the minimal template-based contamination observed does not undermine the validity of our evaluations. Models must effectively utilize learned skills from their training, applying them innovatively to solve new and complex challenges.
Figure A8: Scatter plots illustrating the relationship between official-solution similarity and relative scores for GPT-5 (left) and Grok-4-Fast-Reasoning (right). Both models exhibit weak correlations (Pearson r = 0.136, p = 0.010 for GPT-5; Pearson r = 0.193, p < 0.001 for Grok-4), suggesting minimal impact of template similarity on overall model performance.
Prompt Variants Prompt 1. You are a helpful and knowledgeable AI assistant. Answer the user’s questions clearly and concisely, using correct reasoning and accurate information. Prompt 2. You are an expert IOI competitive programming assistant. For each problem, output only a single complete C++17 solution that compiles and solves the task. Do not include explanations, comments, or any text outside one cpp code block. Read input from standard input and write output to standard output. Prompt 3. You are an expert IOI problem solver. First, think through the solution step by step and explain your reasoning in a few clear paragraphs or bullet points. After that, output a single complete C++17 solution inside one cpp code block that follows your reasoning, reads from standard input, and writes to standard output. Prompt 4. You are a meticulous and safe IOI problem-solving assistant. Always prioritize correctness, handling of corner cases, and numerical stability over constant-factor performance. Carefully verify your reasoning before producing an answer. When writing code, output a single complete C++17 solution in a cpp code block, making sure it compiles, handles invalid or extreme inputs robustly when appropriate, and does not perform unsafe or undefined operations.
E.8 Reasoning Behaviors Analysis